OpenStudy (anonymous):

Find the point on the line y = 3x + 5 that is closest to the origin

OpenStudy (freckles):

calculus?

OpenStudy (freckles):

algebra?

OpenStudy (freckles):

we can do either

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

i can only do calculus

OpenStudy (freckles):

I prefer the calculus way

OpenStudy (freckles):

First step find the distance from (x,3x+5) to (0,0)

OpenStudy (freckles):

You know the distance formula?

OpenStudy (anonymous):

yeah its the distance= sqrt(X2-X1)^2(Y2-Y1)^2)

OpenStudy (anonymous):

there should be a + between the x and y my bad

OpenStudy (freckles):

ok so input the points from our question we are finding the distance from (x,y) to (0,0) where it was given in your question y was 3x+5

OpenStudy (anonymous):

so its distance= sqrt(X2-X)^2+(3x+5-y)^2)

OpenStudy (freckles):

You need to input the points (x,3x+5) and (0,0)

OpenStudy (anonymous):

so sqrt(x^2-(3x+5)^2)

OpenStudy (freckles):

+?

OpenStudy (anonymous):

yeah sorry i hate typing equations

OpenStudy (freckles):

\[d^2=x^2+(3x+5)^2 \\ \text{ let } f(x)=x^2+(3x+5)^2 \] find f' then find the critical numbers

OpenStudy (anonymous):

so f'= 2x +6(3x+5) then i would set that to 0 to find the critical points right?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

so the critical point is X= -20/30

OpenStudy (freckles):

you can reduce that

OpenStudy (anonymous):

right to -2/3

OpenStudy (freckles):

and are you sure that is right?

OpenStudy (freckles):

Did you solve 20x+30=0 for x correctly?

OpenStudy (freckles):

Is not that if ax+b=0 then ax=-b then x=-b/a

OpenStudy (anonymous):

duh i need to pay attention its -3/2

OpenStudy (freckles):

ok so you find the point by replacing x in y=3x+5 to get y

OpenStudy (freckles):

so y=3(-3/2)+5=?

OpenStudy (anonymous):

y=1/2

OpenStudy (freckles):

so you have the point (-3/2,1/2) on y=3x+5 is closest to (0,0) we could have use algebra... y=3x+5 the perpendicular line will have slope -1/3 we want this perpendicular line to go through (0,0) so we have the perpendicular line going through (0,0) is y=-1/3 *x so find the intersection of y=3x+5 and y=-x/3 and we can find the point on y=3x+5 which has the shortest distance from (0,0) so 3x+5=-x/3 multiply 3 on both sides 9x+15=-x 10x+15=0 10x=-15 x=-15/10=-3/2 so y=3x+5=3(-3/2)+5=-9/2+5=-9/2+10/2=1/2 so the point again is (-3/2,1/2)

OpenStudy (anonymous):

thanks for all you help I'm getting most of the problems but everyone once and i while i get tripped in easy algebra thanks again and nice profile pic

OpenStudy (freckles):

lol thanks